# Rings of non-commutative functions and their fields of fractions

### v1

@@ 5 The universal skew field of fractions of $\mathscr{O}_{d}(R)$
The following description of universal skew fields of fractions is taken from [10, Section 7.2]. Given a ring $\mathscr{R}$, a field, $\mathscr{U}$, is said to be the universal skew field of fractions of $\mathscr{R}$ if (i) $\mathscr{R}$ embeds into $\mathscr{U}$ and $\mathscr{U}$ is generated as a skew field by the image of $\mathscr{R}$ under this embedding, and, (ii) $\mathscr{U}$ has the following universal property: If a matrix, $A\in\mathscr{R}^{n\times n}$, becomes invertible under a homomorphism from $\mathscr{R}$, into a skew field, $\mathscr{D}$, then the image of $A$ in $\mathscr{U}^{n\times n}$ is invertible. Equivalently, $\mathscr{U}$ can be characterized using the concept of a local homomorphism. If $\mathscr{E},\mathscr{F}$ are skew fields, a local homomorphism from $\mathscr{E}$ to $\mathscr{F}$ is a ring homomorphism $\varphi:\mathscr{R}_{0}\subseteq\mathscr{E}\rightarrow\mathscr{F}$ whose domain, $\mathscr{R}_{0}\subseteq\mathscr{E}$, is a local subring so that the kernel of $\varphi$ consists precisely of those elements in $\mathscr{R}_{0}$ which are not invertible in $\mathscr{R}_{0}$. A skew field, $\mathscr{U}$, is then the universal skew field of fractions of a ring, $\mathscr{R}$, if (i) $\mathscr{U}$ is generated by the embedding of $\mathscr{R}$ into $\mathscr{U}$, and (ii) every homomorphism from $\mathscr{R}$ into a skew field, $\mathscr{D}$, extends to a local homomorphism from $\mathscr{U}$ into $\mathscr{D}$ whose domain contains $\mathscr{R}$. The universal skew field of fractions of a ring is unique up to isomorphism, if it exists [10, Section 7.2].
@@
    every square full matrix over $\mathscr{R}$ is invertible over $\mathscr{U}$,
see [10, Theorem 7.5.13]. Moreover, [10, Theorem 7.5.13] further asserts that if $\Sigma$ denotes the set of all square full matrices over $\mathscr{R}$, then $\mathscr{U}\simeq\mathscr{R}_{\Sigma}$ is the $\Sigma$-rational closure of $\mathscr{R}$ in $\mathscr{U}$. That is, elements of $\mathscr{U}$ can be identfied, up to a certain equivalence relation given by [10, Lemma 7.4.1], with the entries of the inverses of full matrices over $\mathscr{R}$. It follows, by applying successive Schur complements (and potentially permutations to enable the Schur complements) to any full matrix, $A\in\mathscr{R}^{n\times n}$, that any element of $\mathscr{U}$ can be identified with a certain equivalence class of NC rational expressions composed with elements of $\mathscr{R}$.
Hence, if $f\in\mathscr{M}_{d}(R)$, we can write $f=\mathfrak{r}(h_{1},\cdots,h_{k})\in\mathbb{C}\ \mathclap{\,<}{\left(\right.}\mathfrak{z}_{1},\cdots,\mathfrak{z}_{k}\mathclap{\,\,\,\,\,>}{\left.\right)}\,\,\circ\mathscr{O}_{d}(R)^{1\times k}$ as an NC rational expression composed with elements of $\mathscr{O}_{d}(R)$. In this sense, elements of $\mathscr{M}_{d}(R)$ are “meromorphic NC expressions” in $R\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, and $\mathscr{M}_{d}$ can be thought of as the skew field “globally meromorphic NC expressions”. In particular, any $\mathfrak{r}\in\mathbb{C}\ \mathclap{\,<}{\left(\right.}\mathfrak{z}_{1},\cdots,\mathfrak{z}_{k}\mathclap{\,\,\,\,\,>}{\left.\right)}\,\,$ has a formal realization or formal linear representation of the form $(A,b,c)$, where $A=(A_{0},\cdots,A_{k})$ is a $k+1$-tuple of $m\times m$ complex matrices, and $b,c\in\mathbb{C}^{m}$ [30, Section 4.1][12][22, Section 4.1]. Namely,
@@
$A(\mathfrak{z}):=A_{0}+A_{1}\mathfrak{z}_{1}+\cdots+A_{k}\mathfrak{z}_{k}\in\mathbb{C}\langle\mathfrak{z}_{1},\cdots,\mathfrak{z}_{k}\rangle^{m\times m},$
@@
Since our semifirs, $\mathscr{O}_{d}(R)$, are unital, we have $\mathbb{C}^{m}\subseteq\mathscr{O}_{d}(R)^{m}$ as the constant vector-valued functions. Moreover, any element $f=\mathfrak{r}(h_{1},\cdots,h_{k})\in\mathscr{M}_{d}(R)$, $h_{i}\in\mathscr{O}_{d}(R)$, can then be represented as
@@
In addition to the semifirs, $\mathscr{O}_{d}(R)$, $R\in(0,\infty]$, let $\mathscr{O}_{d}(0)$ denote the semifir of uniformly analytic NC germs at 0 (initially defined and studied in [26]), where $0\in\mathbb{C}^{1\times d}\subseteq\mathbb{C}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ denotes the origin of the NC universe. Namely, a uniformly analytic NC germ at 0 is an equivalence class of uniformly analytic NC functions whose evaluations agree in some uniformly open neighbourhood of 0. If $\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle:=\mathbb{C}\langle\!\langle\mathfrak{z}_{1},\cdots,\mathfrak{z}_{d}\rangle\!\rangle$ denotes the ring of all free formal power series, this is also a semifir by [10, Proposition 2.9.19], and we have the chain of embeddings of semifirs:
@@
for any $r<R\in(0,+\infty]$. An embedding of rings is said to be honest if it preserves fullness of matrices. Each of the embeddings above can be viewed as an embedding into a skew field, e.g. the embedding $\mathscr{O}_{d}(R)\hookrightarrow\mathscr{O}_{d}(r)$ can be viewed as an embedding $\mathscr{O}_{d}(R)\hookrightarrow\mathscr{M}_{d}(r)\supsetneqq\mathscr{O}_{d}(r)$. Hence, if any of these embeddings are honest then they preserve the inner rank of all matrices by [10, Corollary 5.4.10]. By [26, Proposition 5.3], the final embedding, $\mathscr{O}_{d}(0)\hookrightarrow\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle$ is totally inert, hence honest. (The concept, totally inert is a stronger condition that implies honesty). We remark that since $\mathscr{O}_{d}(R)\hookrightarrow\mathscr{O}_{d}(0)$, the evaluations of elements of $\mathscr{O}_{d}(R)$ in $R\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ are generically invertible by [26, Theorem 5.7]. Since elements of $\mathscr{M}_{d}(R)$ can be identified with NC rational expressions in $\mathscr{O}_{d}(R)$, it follows that any $f\in\mathscr{M}_{d}(R)$ which is defined at $0\in\mathbb{C}^{1\times d}$ has well-defined evaluations in a level-wise Zariski-dense NC subset of $R\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$.
However, there could exist, in prinicple, a non-zero element $0\neq f\in\mathscr{M}_{d}(R)$, with $0\in\mathrm{Dom}\,f$, which evaluates identically to 0, everywhere it is defined in $R\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$. That is, such an $f$ would be a “meromorphic identity” in the sense of Amitsur [4]. Here, a celebrated theorem of Amitsur implies that there are no “rational identities”: if $\mathfrak{r}\in\mathbb{C}\ \mathclap{\,<}{\left(\right.}\mathfrak{z}\mathclap{\,\,\,\,\,>}{\left.\right)}\,\,$ is an NC rational function that evaluates to 0 everywhere it is defined in $\mathbb{C}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, then $\mathfrak{r}=0$ [4, Theorem 14] [42, Chapter 8]. Although we expect that there should be an extension of Amitsur’s theorem to $\mathscr{M}_{d}(R)$, $R\in(0,+\infty]$, we have not yet been able to establish this.
@@ 6 Evaluation in stably finite algebras
Recall that $\mathscr{M}_{d}$ denotes the universal skew field of the semifir of uniformly entire NC functions, $\mathscr{O}_{d}$. (In this section, we focus on $\mathscr{M}_{d}=\mathscr{M}_{d}(+\infty)$; our results here extend readily to $\mathscr{M}_{d}(R)$ for any $R>0$.) We will refer to elements of $\mathscr{M}_{d}$ as uniformly meromorphic NC expressions instead of “uniformly meromorphic NC functions” since, as described in the previous section, there may exist elements of $\mathscr{M}_{d}$ that cannot be identified with NC functions. By a theorem of P.M. Cohn, elements of the free skew field, $\mathbb{C}\ \mathclap{\,<}{\left(\right.}\mathfrak{z}\mathclap{\,\,\,\,\,>}{\left.\right)}\,\,$, i.e., NC rational functions, have well-defined evaluations and domains in any stably finite (unital) algebra over $\mathbb{C}$ [11, Theorem 7.3.2], [9]. (Also see [30, Theorem 4.4].) Our goal in this section is to extend this result to evaluation of uniformly meromorphic NC expressions in a suitable class of stably-finite algebras.
@@
is the affine linear pencil of $A$. Setting $\vec{h}:=(h_{1},\cdots,h_{k})$, we will write $A(h_{1}(\mathfrak{z}),\cdots,h_{k}(\mathfrak{z}))=A(\vec{h}(\mathfrak{z}))$. Any element, $f\in\mathscr{M}_{d}$ is then an equivalence class of such rational expressions in $\mathscr{O}_{d}$, where the equivalence relation is described in [10, Lemma 7.4.1]. If $f\in\mathscr{M}_{d}$ is such that there exist $(A,b,c)\in\mathbb{C}^{(n\times n)\cdot(k+1)}\times\mathbb{C}^{n}\times\mathbb{C}^{n}$ and $\vec{h}\in\mathscr{O}_{d}^{1\times k}$ so that the expression given in Equation (6.1) is in the equivalence class of $f$, we will write $f\sim(A,b,c;\vec{h})$ and we will say that $(A,b,c;\vec{h})$ is a formal representation of $f$.
@@
Of course, there is a question whether the value of $f(a)$ is well defined. Namely, it is not obvious whether or not this definition of $f(a)$, for $a\in\mathrm{Dom}_{\mathscr{A}}(f)$, is independent of the choice of formal representation for $f$. That this is indeed always well-defined is the essence of the following theorem.
@@
The following argument is essentially the same as that used in the proof of [30, Theorem 4.4].
@@
$h(\zeta)=\sum_{\omega\in\mathbb{F}}\hat{h}_{\omega}\mathfrak{z}^{\omega};\qquad\qquad 0=R_{h}^{-1}=\lim\sqrt[2n]{\sum_{|\omega|=n}|\hat{h}_{\omega}|^{2}}.$
In order to define evaluation of an entire $h\in\mathscr{O}_{d}$ in a unital, stably-finite algebra, $\mathscr{A}$ over $\mathbb{C}$, we can equip $\mathscr{A}$ with a suitable topology so that this infinite power series converges to an element of $\mathscr{A}$. For example, if $\mathscr{A}$ is a unital, separable and stably-finite $C^{*}$-algebra, then by the Gelfand–Naimark–Segal construction, we can assume, without loss of generality that $\mathscr{A}\subseteq\mathscr{B}(\mathcal{H})$ for a separable and complex Hilbert space, $\mathcal{H}$. In this case given $a_{1},\cdots,a_{d}\in\mathscr{A}\subseteq\mathscr{B}(\mathcal{H})$, the partial sums of the above FPS of $h$, evaluated at $(a_{1},\cdots,a_{d})$ converges absolutely and uniformly on any row-ball in $\mathscr{A}^{1\times d}\subseteq\mathscr{B}(\mathcal{H})^{1\times d}$ by [41, Theorem 1.1]. Namely, for any fixed $r>0$, setting
@@
The proof of [41, Theorem 1.1] explicitly uses the $C^{*}$-identity for operators. Namely, given $A\in\mathscr{B}(\mathcal{H},\mathcal{J})$, where $\mathcal{H},\mathcal{J}$ are complex Hilbert spaces, $\|A^{*}A\|=\|A\|^{2}$. Explicitly, if $h(\mathfrak{z})=\sum_{\omega\in\mathbb{W}_{d}}\hat{h}_{\omega}\mathfrak{z}^{\omega}\in\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle$ has radius of convergence $R_{h}>0$, and $A\in\mathscr{B}(\mathcal{H})^{1\times d}$, then one estimates:
@@
$=$ $\sum_{\ell=0}^{\infty}\left\|\mathrm{row}(A^{\omega})_{|\omega|=\ell}\mathrm{col}(\hat{h}_{\omega})_{|\omega|=\ell}\right\|$
@@
Given a $d$-tuple of bounded linear operators, $A\in\mathscr{B}(\mathcal{H})^{1\times d}$ one defines the joint spectral radius of $A$, $\rho_{\sigma}(A)$ as the square root of the spectral radius of the bounded linear map, $\mathrm{Ad}_{A,A^{*}}$, viewed as an element of the unital Banach algebra $\mathscr{B}(X)$, where $X:=\mathscr{B}(\mathcal{H})$ [39]. Namely,
@@
so that “multiplication is completely contractive” and $(A_{\rho},\|\cdot\|_{\rho})$ is an abstract operator algebra in sense of Blecher, Ruan and Sinclair [33, Corollary 16.7]. We will further assume that our matrix semi-norms are ‘unital’ in the sense that $\rho(1)=1$. In this case $(A_{\rho},\|\cdot\|_{\rho})$ is a unital abstract operator algebra, and hence is completely isometrically isomorphic to a unital algebra of operators on a Hilbert space by the Blecher–Ruan–Sinclair theorem.
@@
A (unital) local operator algebra is a unital, separable, Hausdorff and complete topological algebra, $\mathscr{A}$, over $\mathbb{C}$, equipped with a family of unital and matrix-multiplicative semi-norms that generate the topology on $\mathscr{A}$.
The inverse limit of any family of unital operator algebras will be a local operator algebra. For example, each of the ball semifirs of uniformly analytic NC functions constructed in this paper, $\mathscr{O}_{d}(R)$, $R>0$, are local operator algebras. Conversely, any local operator algebra can be viewed as an inverse limit of abstract operator algebras [53]. If $\mathscr{A}$ is a local operator algebra, $a=(a_{1},\cdots,a_{d})\in\mathscr{A}^{1\times d}$ and $h\in\mathscr{O}_{d}$, the free polynomial partial sums, $h_{n}(a)$ then converge to an element, $h(a)$, in $\mathscr{A}$. Indeed, to prove this, one need only show that given any $\rho\in\Lambda$, that the sequence $h_{n}(a+N_{\rho})$ is Cauchy in the abstract, unital and separable operator algebra $A/N_{\rho}$, and this follows immediately from [41, Theorem 1.1], since $h$ has infinite Cauchy–Hadamard radius of convergence and we can view $A/N_{\rho}$ as an algebra of operators on a separable, complex Hilbert space. That is, if $\mathscr{A}$ is a unital local operator algebra, then for any $d$-tuple $a=(a_{1},\cdots,a_{d})\in\mathscr{A}^{1\times d}$, the unital homomorphism, $\varphi_{a}:\mathscr{O}_{d}\rightarrow\mathscr{A}$ is well-defined. Hence, if a unital local operator algebra is also stably-finite, Theorem 6.1 applies and shows that we have well-defined $\mathscr{A}$-domains and evaluations for any $f\in\mathscr{M}_{d}$.

### v2

@@ 5 The universal skew field of fractions of $\mathscr{O}_{d}(R)$
The following description of universal skew fields of fractions is taken from [10, Section 7.2]. Given a ring $\mathscr{R}$, a field, $\mathscr{U}$, is said to be the universal skew field of fractions of $\mathscr{R}$ if (i) $\mathscr{R}$ embeds into $\mathscr{U}$ and $\mathscr{U}$ is generated as a skew field by the image of $\mathscr{R}$ under this embedding, and, (ii) $\mathscr{U}$ has the following universal property: If a matrix, $A\in\mathscr{R}^{n\times n}$, becomes invertible under a homomorphism from $\mathscr{R}$, into a skew field, $\mathscr{D}$, then the image of $A$ in $\mathscr{U}^{n\times n}$ is invertible. Equivalently, $\mathscr{U}$ can be characterized using the concept of a local homomorphism. If $\mathscr{E},\mathscr{F}$ are skew fields, a local homomorphism from $\mathscr{E}$ to $\mathscr{F}$ is a ring homomorphism $\varphi:\mathscr{R}_{0}\subseteq\mathscr{E}\rightarrow\mathscr{F}$ whose domain, $\mathscr{R}_{0}\subseteq\mathscr{E}$, is a local subring so that the kernel of $\varphi$ consists precisely of those elements in $\mathscr{R}_{0}$ that are not invertible in $\mathscr{R}_{0}$. A skew field, $\mathscr{U}$, is then the universal skew field of fractions of a ring, $\mathscr{R}$, if (i) $\mathscr{U}$ is generated by the embedding of $\mathscr{R}$ into $\mathscr{U}$, and (ii) every homomorphism from $\mathscr{R}$ into a skew field, $\mathscr{D}$, extends to a local homomorphism from $\mathscr{U}$ into $\mathscr{D}$ whose domain contains $\mathscr{R}$. The universal skew field of fractions of a ring is unique up to isomorphism, if it exists [10, Section 7.2].
@@
    every full, square matrix over $\mathscr{R}$ is invertible over $\mathscr{U}$,
see [10, Theorem 7.5.13]. Moreover, [10, Theorem 7.5.13] further asserts that if $\Sigma$ denotes the set of all square full matrices over $\mathscr{R}$, then $\mathscr{U}\simeq\mathscr{R}_{\Sigma}$ is the $\Sigma$-rational closure of $\mathscr{R}$ in $\mathscr{U}$. That is, elements of $\mathscr{U}$ can be identfied, up to a certain equivalence relation given by [10, Section 7.4, Lemma 7.4.1], with the entries of the inverses of full matrices over $\mathscr{R}$. It follows, by applying successive Schur complements (and potentially permutations to enable the Schur complements) to any full matrix, $A\in\mathscr{R}^{n\times n}$, that any element of $\mathscr{U}$ can be identified with a certain equivalence class of NC rational expressions composed with elements of $\mathscr{R}$.
Hence, if $f\in\mathscr{M}_{d}(R)$, we can write $f=\mathfrak{r}(h_{1},\cdots,h_{k})\in\mathbb{C}\ \mathclap{\,<}{\left(\right.}\mathfrak{z}_{1},\cdots,\mathfrak{z}_{k}\mathclap{\,\,\,\,\,>}{\left.\right)}\,\,\circ\mathscr{O}_{d}(R)^{1\times k}$ as an NC rational expression composed with elements of $\mathscr{O}_{d}(R)$. In this sense, elements of $\mathscr{M}_{d}(R)$ are “meromorphic NC expressions” in $R\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, and $\mathscr{M}_{d}$ can be thought of as the skew field “globally meromorphic NC expressions”. In particular, any $\mathfrak{r}\in\mathbb{C}\ \mathclap{\,<}{\left(\right.}\mathfrak{z}_{1},\cdots,\mathfrak{z}_{k}\mathclap{\,\,\,\,\,>}{\left.\right)}\,\,$ has a formal realization or formal linear representation of the form $(A,b,c)$, where $A=(A_{0},\cdots,A_{k})$ is a $k+1$-tuple of $m\times m$ complex matrices, and $b,c\in\mathbb{C}^{m}$ [31, Section 4.1][12][23, Section 4.1]. Namely,
@@
$A(\mathfrak{z}):=A_{0}\cdot 1+A_{1}\mathfrak{z}_{1}+\cdots+A_{k}\mathfrak{z}_{k}\in\mathbb{C}\langle\mathfrak{z}_{1},\cdots,\mathfrak{z}_{k}\rangle^{m\times m},$
@@
Since our semifirs, $\mathscr{O}_{d}(R)$, are unital, we have $\mathbb{C}^{m}\subseteq\mathscr{O}_{d}(R)^{m}$ as the constant vector-valued functions. Any element $f=\mathfrak{r}(h_{1},\cdots,h_{k})\in\mathscr{M}_{d}(R)$, $h_{i}\in\mathscr{O}_{d}(R)$, can then be represented as
@@
In addition to the semifirs, $\mathscr{O}_{d}(R)$, $R\in(0,\infty]$, let $\mathscr{O}_{d}(0)$ denote the semifir of uniformly analytic NC germs at 0 (initially defined and studied in [27]), where $0\in\mathbb{C}^{1\times d}\subseteq\mathbb{C}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ denotes the origin of the NC universe. Namely, a uniformly analytic NC germ at 0 is an equivalence class of uniformly analytic NC functions whose evaluations agree in some uniformly open neighbourhood of 0. If $\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle=\mathbb{C}\langle\!\langle\mathfrak{z}_{1},\cdots,\mathfrak{z}_{d}\rangle\!\rangle$ denotes the ring of all free formal power series, this is also a semifir by [10, Proposition 2.9.19], and we have the chain of embeddings of semifirs:
@@
for any $r<R\in(0,+\infty]$. An embedding of rings is said to be honest if it preserves fullness of matrices. Each of the embeddings above can be viewed as an embedding into a skew field, e.g. the embedding $\mathscr{O}_{d}(R)\hookrightarrow\mathscr{O}_{d}(r)$ can be viewed as an embedding $\mathscr{O}_{d}(R)\hookrightarrow\mathscr{M}_{d}(r)\supsetneqq\mathscr{O}_{d}(r)$. Hence, if any of these embeddings are honest then they preserve the inner rank of all matrices by [10, Corollary 5.4.10]. By [27, Proposition 5.3], the final embedding, $\mathscr{O}_{d}(0)\hookrightarrow\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle$ is totally inert, hence honest. (The concept, totally inert is a stronger condition that implies honesty.) We remark that since $\mathscr{O}_{d}(R)\hookrightarrow\mathscr{O}_{d}(0)$, the evaluations of elements of $\mathscr{O}_{d}(R)$ in $R\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$ are generically invertible by [27, Theorem 5.7]. Since elements of $\mathscr{M}_{d}(R)$ can be identified with NC rational expressions in $\mathscr{O}_{d}(R)$, it follows that any $f\in\mathscr{M}_{d}(R)$ which is defined at $0\in\mathbb{C}^{1\times d}$ has well-defined evaluations in a level-wise Zariski-dense NC subset of $R\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$.
However, there could exist, in prinicple, a non-zero element $0\neq f\in\mathscr{M}_{d}(R)$, with $0\in\mathrm{Dom}\,f$, which evaluates identically to 0, everywhere it is defined in $R\cdot\mathbb{B}^{(\mathbb{N}\times\mathbb{N})\cdot d}$. That is, such an $f$ would be a “meromorphic identity” in the sense of Amitsur [4]. Here, a celebrated theorem of Amitsur implies that there are no “rational identities”: If $\mathfrak{r}\in\mathbb{C}\ \mathclap{\,<}{\left(\right.}\mathfrak{z}\mathclap{\,\,\,\,\,>}{\left.\right)}\,\,$ is an NC rational function that evaluates to 0 everywhere it is defined in $\mathbb{C}^{(\mathbb{N}\times\mathbb{N})\cdot d}$, then $\mathfrak{r}=0$ [4, Theorem 14] [44, Chapter 8]. Although we expect that there should be an extension of Amitsur’s theorem to $\mathscr{M}_{d}(R)$, $R\in(0,+\infty]$, we have not yet been able to establish this.
@@ 6 Evaluation in stably finite algebras
Recall that $\mathscr{M}_{d}$ denotes the universal skew field of the semifir of uniformly entire NC functions, $\mathscr{O}_{d}$. (In this section, we focus on $\mathscr{M}_{d}=\mathscr{M}_{d}(+\infty)$; our results here extend readily to $\mathscr{M}_{d}(R)$ for any $R>0$.) We will refer to elements of $\mathscr{M}_{d}$ as uniformly meromorphic NC expressions instead of “uniformly meromorphic NC functions” since, as described in the previous section, there may exist elements of $\mathscr{M}_{d}$ that cannot be identified with NC functions. By a theorem of P.M. Cohn, elements of the free skew field, $\mathbb{C}\ \mathclap{\,<}{\left(\right.}\mathfrak{z}\mathclap{\,\,\,\,\,>}{\left.\right)}\,\,$, i.e., NC rational functions, have well-defined evaluations and domains in any stably finite (unital) algebra over $\mathbb{C}$ [11, Theorem 7.3.2], [9]. (Also see [31, Theorem 4.4].) Our goal in this section is to extend this result to evaluation of uniformly meromorphic NC expressions in a suitable class of stably-finite algebras.
@@
is the affine linear pencil of $A$. Setting $\vec{h}:=(h_{1},\cdots,h_{k})$, we will write $A(h_{1}(\mathfrak{z}),\cdots,h_{k}(\mathfrak{z}))=A(\vec{h}(\mathfrak{z}))$. Any element, $f\in\mathscr{M}_{d}$ is then an equivalence class of such rational expressions in $\mathscr{O}_{d}$, where the equivalence relation is described in [10, Lemma 7.4.1]. If $f\in\mathscr{M}_{d}$ is such that there exists a $k\in\mathbb{N}$ and $(A,b,c)\in\mathbb{C}^{(n\times n)\cdot(k+1)}\times\mathbb{C}^{n}\times\mathbb{C}^{n}$ and $\vec{h}\in\mathscr{O}_{d}^{1\times k}$ so that the expression given in Equation (6.1) is in the equivalence class of $f$, we will write $f\sim(A,b,c;\vec{h})$ and we will say that $(A,b,c;\vec{h})$ is a formal representation of $f$.
@@
Of course, there is the question of whether the value of $f(a)$ is well defined. Namely, it is not obvious whether or not this definition of $f(a)$, for $a\in\mathrm{Dom}_{\mathscr{A}}(f)$, is independent of the choice of formal representation for $f$. That this is indeed always well-defined is the essence of the following theorem.
@@
The following argument is essentially the same as that used in the proof of [31, Theorem 4.4].
@@
$h(\zeta)=\sum_{\omega\in\mathbb{W}_{d}}\hat{h}_{\omega}\mathfrak{z}^{\omega};\qquad\qquad 0=\frac{1}{R_{h}}=\lim_{n\rightarrow\infty}\sqrt[2n]{\sum_{|\omega|=n}|\hat{h}_{\omega}|^{2}}.$
In order to define evaluation of an entire $h\in\mathscr{O}_{d}$ in a unital, stably-finite algebra, $\mathscr{A}$ over $\mathbb{C}$, we can equip $\mathscr{A}$ with a suitable topology so that this infinite power series converges to an element of $\mathscr{A}$. For example, if $\mathscr{A}$ is a unital, separable and stably-finite $C^{*}$-algebra, then by the Gelfand–Naimark–Segal construction, we can assume, without loss of generality that $\mathscr{A}\subseteq\mathscr{B}(\mathcal{H})$ for a separable and complex Hilbert space, $\mathcal{H}$. In this case given $a_{1},\cdots,a_{d}\in\mathscr{A}\subseteq\mathscr{B}(\mathcal{H})$, the partial sums of the above FPS of $h$, evaluated at $(a_{1},\cdots,a_{d})$ converge absolutely and uniformly on any row-ball in $\mathscr{A}^{1\times d}\subseteq\mathscr{B}(\mathcal{H})^{1\times d}$ by [43, Theorem 1.1]. Namely, for any fixed $r>0$, setting
@@
The proof of [43, Theorem 1.1] explicitly uses the $C^{*}$-identity for operators. Namely, given $A\in\mathscr{B}(\mathcal{H},\mathcal{J})$, where $\mathcal{H},\mathcal{J}$ are complex Hilbert spaces, $\|A^{*}A\|=\|A\|^{2}$. Explicitly, if $h(\mathfrak{z})=\sum_{\omega\in\mathbb{W}_{d}}\hat{h}_{\omega}\mathfrak{z}^{\omega}\in\mathbb{C}\langle\!\langle\mathfrak{z}\rangle\!\rangle$ has radius of convergence $R_{h}>0$, and $A\in\mathscr{B}(\mathcal{H})^{1\times d}$, then one estimates:
@@
$=$ $\sum_{\ell=0}^{\infty}\left\|\mathrm{row}(A^{\omega})_{|\omega|=\ell}\cdot\mathrm{col}(\hat{h}_{\omega})_{|\omega|=\ell}\right\|$
@@
Given a $d$-tuple of bounded linear operators, $A\in\mathscr{B}(\mathcal{H})^{1\times d}$ one defines the joint spectral radius of $A$, $\rho_{\sigma}(A)$ as the square root of the spectral radius of the bounded linear map, $\mathrm{Ad}_{A,A^{*}}$, viewed as an element of the unital Banach algebra $\mathscr{B}(X)$, where $X=\mathscr{B}(\mathcal{H})$ [41]. Namely,
@@
so that “multiplication is completely contractive” and $(A_{\rho},\|\cdot\|_{\rho})$ is an abstract operator algebra in sense of Blecher, Ruan and Sinclair [34, Corollary 16.7]. We will further assume that our matrix semi-norms are ‘unital’ in the sense that $\rho(1)=1$. In this case $(A_{\rho},\|\cdot\|_{\rho})$ is a unital abstract operator algebra, and hence is completely isometrically isomorphic to a unital algebra of operators on a Hilbert space by the Blecher–Ruan–Sinclair theorem.
@@
A (unital) local operator algebra is a unital and separable topological algebra, $\mathscr{A}$, over $\mathbb{C}$, equipped with a family of unital and matrix-multiplicative semi-norms that generate the topology on $\mathscr{A}$.
The inverse limit of any family of unital operator algebras will be a local operator algebra. For example, each of the ball semifirs of uniformly analytic NC functions constructed in this paper, $\mathscr{O}_{d}(R)$, $R>0$, are local operator algebras. Conversely, any local operator algebra can be viewed as an inverse limit of abstract operator algebras [55]. If $\mathscr{A}$ is a local operator algebra, $a=(a_{1},\cdots,a_{d})\in\mathscr{A}^{1\times d}$ and $h\in\mathscr{O}_{d}$, the free polynomial partial sums, $h_{n}(a)$ then converge to an element, $h(a)$, in $\mathscr{A}$. Indeed, to prove this, one need only show that given any $\rho\in\Lambda$, that the sequence $h_{n}(a+N_{\rho})$ is Cauchy in the abstract, unital and separable operator algebra $A/N_{\rho}$, and this follows immediately from [43, Theorem 1.1], since $h$ has infinite Cauchy–Hadamard radius of convergence and we can view $A/N_{\rho}$ as an algebra of operators on a separable, complex Hilbert space. That is, if $\mathscr{A}$ is a unital local operator algebra, then for any $d$-tuple $a=(a_{1},\cdots,a_{d})\in\mathscr{A}^{1\times d}$, the unital homomorphism, $\varphi_{a}:\mathscr{O}_{d}\rightarrow\mathscr{A}$ is well-defined. Hence, if a unital local operator algebra is also stably-finite, Theorem 6.1 applies and shows that we have well-defined $\mathscr{A}$-domains and evaluations for any $f\in\mathscr{M}_{d}$. That is, we have shown the following.
Corollary 6.4.
Let $\mathscr{A}$ be a unital local operator algebra. Then for any $a=(a_{1},\cdots,a_{d})\in\mathscr{A}^{1\times d}$, there is a well-defined unital homomorphism $\varphi_{a}:\mathscr{O}_{d}\rightarrow\mathscr{A}$ obeying $\varphi_{a}(\mathfrak{z}_{i})=a_{i}$. If $\mathscr{A}$ is also stably-finite, then any $f\in\mathscr{M}_{d}$ has a well-defined $\mathscr{A}$-domain and evaluations,
$\mathrm{Dom}_{\mathscr{A}}(f):=\{a\in\mathscr{A}^{1\times d}|\ f\sim(A,b,c;\vec{h})\ \mbox{and}\ A(\vec{h}(a))^{-1}\,\exists\},$
$f(a)=1_{\mathscr{A}}\otimes b^{*}A(\vec{h}(a))^{-1}1_{\mathscr{A}}\otimes c;\qquad\qquad a\in\mathrm{Dom}_{\mathscr{A}}(f).$
In particular, each $\mathscr{O}_{d}(R)$, $R\in(0,+\infty]$ is a stably-finite local operator algebra.